### v2

@@ 4.2 Experimental Setting
Proposed method (ResNet-50 as backbone) 95.11% 95.81% 94.04% 94.92%
Proposed method (ResNet-101 as backbone) 94.18% 94.76% 92.79% 93.76%
@@
Reshan et al. (MobileNet as backbone) reshan2023detection 0.85% 91.41% 95.28% 91.41%
Proposed method (ResNet-50 as backbone) 91.03% 91.54% 94.36% 92.93%
Proposed method (ResNet-101 as backbone) 92.79% 93.45% 95.13% 94.28%
@@
Backbones Results of proposed method Acuracy Precision Recall F1-score AUC
ResNet-50 on original images 91.23% 90.94% 85.60% 88.19% 89.41%
on predicted masks 95.11% 95.81% 94.04% 94.92% 95.07%
ReNet101 on original images 90.22% 88.16% 87.04% 87.6% 90.16%
on predicted masks 94.18% 94.76% 92.79% 93.76% 94.28%
4.3 Evaluation Metrics
Accuracy=True Positives+True NegativesTotal Number of $\text{Accuracy}=\frac{\text{True Positives}+\text{True Negatives}}{\text{TotalNumber of Instance}}$ (10)
Precision=True PositivesTrue Positives+True $\text{Precision}=\frac{\text{True Positives}}{\text{True Positives}+\text{TrueNegatives}}$ (11)
Recall=True PositivesTrue Positives+False $\text{Recall}=\frac{\text{True Positives}}{\text{True Positives}+\text{False Negatives}}$ (12)
[…]
@@ 4.3 Evaluation Metrics
4.4 Comparison with State-of-the-Art
We compared our proposed method with several state-of-the-art methods on both datasets. As shown in Table 1 and Table 2, our proposed method achieved the best results across all evaluation metrics, including accuracy, precision, recall, and F1 score, using both ResNet-50 and ResNet-101 backbones. This demonstrates the effectiveness of our approach in classifying pneumonia from chest X-ray images.
Our method leverages the strengths of pre-trained ResNet models to extract robust feature representations. By freezing the pre-trained layers, we focus the learning on the newly added layers, reducing the number of learnable parameters to only 0.27 million.
@@ 4.5 Ablation study
The results of the ablation study are summarized in Table 3. The table displays classification metrics, including accuracy, precision, recall, F1-score, and AUC, for both ResNet-50 and ResNet-101 backbones under the two different input scenarios.
@@
X X 87.73% 80.55% 79.62% 80.08%
X X 87.73% 80.63% 80.47% 80.55%
X X X 91.23% 90.94% 85.60% 88.19%
ResNet-101 X 83.93% 74.21% 73.92% 74.06%
X X 85.56% 78.21% 80.4% 79.29%
X X 85.71% 78.59% 81.11% 79.83%
X X X 90.22% 88.16% 87.04% 87.6%
The results clearly demonstrate the benefit of incorporating segmentation masks in the classification process. For both ResNet-50 and ResNet-101 backbones, the model achieves higher accuracy, precision, recall, F1-score, and AUC when trained on images with predicted lung masks compared to the raw images. Specifically, the accuracy improves by 2 percentage points for ResNet-50 and by 2.73 percentage points for ResNet-101. Similarly, the precision, recall, F1-score, and AUC all show substantial improvements.
@@
For the ResNet-50 backbone, the confusion matrix in Figure 3(a) shows the results on original images. The model correctly classifies 88 COVID-19 cases, 271 normal cases, and 816 pneumonia cases, with a notable number of misclassifications, particularly in the COVID-19 and pneumonia categories. When using predicted masks, as shown in Figure 3(b), the model’s performance improves significantly, correctly classifying 112 COVID-19 cases, 279 normal cases, and 834 pneumonia cases. The number of misclassifications decreases across all categories, highlighting the benefit of segmentation in isolating relevant features.
For the ResNet-101 backbone, the confusion matrix in Figure 3(c) displays the results on original images, where the model correctly classifies 94 COVID-19 cases, 278 normal cases, and 790 pneumonia cases. However, the misclassifications are more pronounced compared to ResNet-50, particularly in the pneumonia category. With the predicted masks, as shown in Figure 3(d), the performance improves, with correct classifications of 107 COVID-19 cases, 286 normal cases, and 820 pneumonia cases. This reduction in misclassifications further supports the effectiveness of incorporating segmentation masks.
@@
The findings reveal a significant enhancement in model performance when both multi-scale feature maps and the transformer are employed alongside the baseline configuration. For instance, with the ResNet-50 backbone, the accuracy improves from 84.62% (baseline) to 91.23% when all components are utilized. Similarly, the ResNet-101 backbone exhibits a notable increase in accuracy from 83.93% to 90.22%. These results underscore the effectiveness of our proposed innovations, illustrating that the integration of multi-scale feature maps and transformer elements not only enhances overall accuracy but also boosts precision, recall, and F1-score, which are crucial for the reliability of classification tasks. This highlights the importance of these key components in achieving improved performance in deep learning models for image analysis.
@@ 5 Conclusion
The high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset confirm the robustness and reliability of our model. The reduction in the number of parameters compared to other state-of-the-art transformer models highlights our contribution to creating a more efficient yet powerful diagnostic tool suitable for deployment in resource-constrained environments. Our work paves the way for future research in several areas. Future work could explore further optimization of the transformer module to enhance performance and reduce computational complexity. Additionally, expanding the dataset to include a broader variety of pneumonia cases and other respiratory diseases could improve the model’s generalization.

### v3

@@ 4.4 Comparison with State-of-the-Art
Zhao et al. (Channel-Attention Capsule) zhao2023dcacorrcapsnet 90.43% 90.81% 90.43% 90.40%
CNN-based van2024large 92.52% - - -
CNN-based van2024large 91.05% - - -
Proposed method (ResNet-50 as backbone) 96.04%$\pm$0.5 96.70%$\pm$0.5 94.90%$\pm$0.6 95.77%$\pm$0.5
Proposed method (ResNet-101 as backbone) 95.19%$\pm$0.6 95.59%$\pm$0.5 94.19%$\pm$0.6 94.86%$\pm$0.6
@@
Reshan et al. (MobileNet as backbone) reshan2023detection 90.85% 91.41% 95.28% 91.41%
Reshan et al. (ResNet152B2 as backbone) reshan2023detection 84.65% 82.38% 99.21% 90.02%
Reshan et al. (DenseNet121 as backbone) reshan2023detection 88.90% 88.33% 96.87% 92.41%
Reshan et al. (Xception as backbone) reshan2023detection 87.59% 91.75% 90.32% 91.03%
Reshan et al. (EfficientNet as backbone) reshan2023detection 51.02% 86.21% 45.85% 90.10%
Jiang et al. (MP-ViT) jiang2022multisemantic 91.19% 91.82% 89.36% 90.34%
ViT in mabrouk2022pneumonia 92.45% 92.47% 92.44% 92.47%
Proposed method (ResNet-50 as backbone) 91.67%$\pm$0.6 92.04%$\pm$0.5 94.87%$\pm$0.5 93.43%$\pm$0.6
Proposed method (ResNet-101 as backbone) 93.75%$\pm$0.5 93.98%$\pm$0.5 96.16%$\pm$0.5 95.05%$\pm$0.5
@@
Backbones Results of proposed method Acuracy Precision Recall F1-score
ResNet-50 on original images 91.23% 90.94% 85.60% 88.19%
on predicted masks 96.04% 96.70% 94.90% 95.77%
ReNet101 on original images 90.22% 88.16% 87.04% 87.6%
on predicted masks 95.19% 95.59% 94.19% 94.86%
4.4.1 Segmentation
Our lightweight transformer-enhanced TransUNet, initially designed to improve the accuracy of pneumonia classification by providing precise lung segmentation, also demonstrates superior performance in the segmentation task itself. While the primary goal of our segmentation model was to isolate lung regions for better classification, as shown in Table 1, its ability to outperform state-of-the-art segmentation methods highlights the effectiveness of our design choices. By integrating transformer-based attention mechanisms into the U-Net architecture, our model captures both local fine-grained details and global contextual information, which are essential for accurate lung segmentation. To ensure the robustness and reliability of our results, we trained and tested the model three times, with consistent performance across all runs. This dual capability—enhancing both segmentation and classification—sets our approach apart from existing methods, which often focus on one task at the expense of the other.
4.4.2 Classification
We evaluated the performance of our proposed method on the ”Cohen” and ”Kermany” datasets, comparing it against several state-of-the-art methods. As shown in Table 2 and Table 3, our approach outperformed existing methods across all evaluation metrics, including accuracy, precision, recall, and F1-score, using both ResNet-50 and ResNet-101 backbones. On the ”Cohen” dataset, our method achieved an accuracy of 96.04% with ResNet-50 and 95.19% with ResNet-101, surpassing the best-performing existing method (Zhao et al. zhao2023dcacorrcapsnet ) by a significant margin. Similarly, on the ”Kermany” dataset, our method achieved an accuracy of 93.75% with ResNet-101 and 91.67% with ResNet-50, outperforming recent transformer-based models like MP-ViT jiang2022multisemantic and ViT mabrouk2022pneumonia .
The superior performance of our method can be attributed to several key design choices. First, by leveraging pre-trained ResNet models, we extract robust multi-scale feature representations that capture both low-level and high-level patterns in chest X-rays. Freezing the pre-trained layers allows the model to focus on learning task-specific features in the newly added layers, reducing overfitting and improving generalization. Second, our lightweight transformer module enhances the model’s ability to capture global contextual information, which is critical for distinguishing subtle pneumonia-related patterns from complex backgrounds. This integration of local and global feature extraction ensures robust performance across diverse datasets.
Furthermore, our method demonstrates consistent performance across both datasets, highlighting its generalizability and adaptability to different clinical settings. For instance, on the ”Cohen” dataset, our method achieved an F1-score of 95.77%, significantly higher than the 83.48% reported by Gadza et al. gazda2021self . Similarly, on the ”Kermany” dataset, our method achieved an F1-score of 95.05%, outperforming the 92.41% achieved by Reshan et al. reshan2023detection using DenseNet121. These results underscore the effectiveness of our approach in addressing the challenges of pneumonia classification, such as class imbalance, subtle visual indicators, and domain shifts between datasets.
In summary, our proposed method not only achieves state-of-the-art performance but also demonstrates computational efficiency and generalizability, making it a promising solution for real-world clinical applications. The integration of pre-trained ResNet models with a lightweight transformer module provides a robust framework for accurate and efficient pneumonia detection, addressing the limitations of existing methods.
@@ 4.5 Ablation study
The results of the ablation study are summarized in Table 4. The table displays classification metrics, including accuracy, precision, recall, and F1-score, for both ResNet-50 and ResNet-101 backbones under the two different input scenarios.
@@
Training time
(per image)
Learnable
parameters
ResNet-50 X 84.62% 75.68% 75.10% 75.39% 0.006s 0.65 M
X X 87.73% 80.55% 79.62% 80.08% 0.009s 0.85 M
X X 87.73% 80.63% 80.47% 80.55% 0.012s 1.15 M
X X X 91.23% 90.94% 85.60% 88.19% 0.014s 1.97 M
ResNet-101 X 83.93% 74.21% 73.92% 74.06% 0.007s 0.65 M
X X 85.56% 78.21% 80.4% 79.29% 0.011s 0.85 M
X X 85.71% 78.59% 81.11% 79.83% 0.016s 1.15 M
X X X 90.22% 88.16% 87.04% 87.6% 0.024s 1.97 M
Table 6: The effect of using different ResNet blocks on ”Cohen” dataset
Backbones Block 2 Block 3 Block 4 Accuracy Precision Recall F1-score
@@
X X 85.86% 77.03% 74.37% 75.68%
X X 86.64% 79.36% 76.15% 77.72%
X X X 86.72% 79.02% 76.48% 77.73%
Combined X 87.73% 80.55% 79.62% 80.08%
The results clearly demonstrate the benefit of incorporating segmentation masks in the classification process. For both ResNet-50 and ResNet-101 backbones, the model achieves higher accuracy, precision, recall, and F1-score when trained on images with predicted lung masks compared to the raw images. Specifically, the accuracy improves by almost five percentage points for both ResNet-50 and ResNet-101. Similarly, the precision, recall, and F1-score all show substantial improvements.
@@
For the ResNet-50 backbone, the confusion matrix in Figure 4(a) shows the results on original images. The model correctly classifies 88 COVID-19 cases, 271 normal cases, and 816 pneumonia cases, with a notable number of misclassifications, particularly in the COVID-19 and pneumonia categories. When using predicted masks, as shown in Figure 4(b), the model’s performance improves significantly, correctly classifying 112 COVID-19 cases, 285 normal cases, and 840 pneumonia cases. The number of misclassifications decreases across all categories, highlighting the benefit of segmentation in isolating relevant features.
For the ResNet-101 backbone, the confusion matrix in Figure 4(c) displays the results on original images, where the model correctly classifies 94 COVID-19 cases, 278 normal cases, and 790 pneumonia cases. However, the misclassifications are more pronounced compared to ResNet-50, particularly in the pneumonia category. With the predicted masks, as shown in Figure 4(d), the performance improves, with correct classifications of 109 COVID-19 cases, 292 normal cases, and 825 pneumonia cases. This reduction in misclassifications further supports the effectiveness of incorporating segmentation masks.
@@
The findings reveal a significant enhancement in model performance when both multi-scale feature maps and the transformer are employed alongside the baseline configuration. For instance, with the ResNet-50 backbone, the accuracy improves from 84.62% (baseline) to 91.23% when all components are utilized. Similarly, the ResNet-101 backbone exhibits a notable increase in accuracy from 83.93% to 90.22%. These results underscore the effectiveness of our proposed innovations, illustrating that the integration of multi-scale feature maps and transformer elements not only enhances overall accuracy but also boosts precision, recall, and F1-score, which are crucial for the reliability of classification tasks. This highlights the importance of these key components in achieving improved performance in deep learning models for image analysis. Additionally, the inclusion of training time (per image) and learnable parameters underscores the computational efficiency and scalability of our approach, ensuring its suitability for resource-constrained environments.
Additionally, we conducted an ablation study to evaluate the contribution of each ResNet block (Blocks 2, 3, and 4) to the classification performance. The results, as shown in Table 6, demonstrate that intermediate layers (Blocks 2 and 3) play a significant role in improving performance, but their contribution is not as critical as that of Block 4. For instance, using only Block 4 achieves an accuracy of 84.62%, while adding Block 3 (without Block 2) improves the accuracy to 86.64%. This indicates that mid-level features from Block 3 provide additional discriminative information, enhancing the model’s ability to detect subtle patterns in chest X-rays. However, the inclusion of Block 2, which captures low-level features, results in only a marginal improvement in accuracy (86.72%), suggesting that while low-level features are beneficial, they are less impactful compared to mid- and high-level features.
Interestingly, the best performance is achieved when using two feature maps: Block 4 ($B^{4}$) and the merged feature map ($B^{merged}$), which combines Blocks 2 and 3. This configuration achieves an accuracy of 87.73%, a precision of 80.55%, and an F1-score of 80.08%, outperforming the scenario where all three blocks are used independently. This result highlights the importance of separating high-level features (Block 4) from intermediate features (Blocks 2 and 3). While intermediate layers provide valuable contextual information, they are not as discriminative as the high-level features from Block 4. By merging Blocks 2 and 3 into a single feature map ($B^{merged}$), we reduce redundancy and computational complexity while preserving the benefits of multi-scale feature extraction. This design choice ensures that the model focuses on the most relevant features for pneumonia detection, leading to improved performance and efficiency.
@@ 5 Conclusion
The high accuracy rates of 93.75% on the ”Kermany” dataset and 96.04% on the ”Cohen” dataset confirm the robustness and reliability of our model. The reduction in the number of parameters compared to other state-of-the-art transformer models highlights our contribution to creating a more efficient yet powerful diagnostic tool suitable for deployment in resource-constrained environments. Our work paves the way for future research in several areas. Future work could explore further optimization of the transformer module to enhance performance and reduce computational complexity. Additionally, expanding the dataset to include a broader variety of pneumonia cases and other respiratory diseases could improve the model’s generalization.